OpenStudy (anonymous):

Find the equation of the line (-4, 6) and perpendicular to y = -3x + 4

OpenStudy (anonymous):

For prependicular slope we do the opposite reciprocal of the slope we already have. This just means, we do the opposite of -3, which is 1/3 and that's our slope for this prependicular line. Now we use the point-slope formula to find our equation. y-y1=m(x-x1) [All we do is put in the set of coords. here] y-6=1/3(x+4) y-6=1/3x+4/3 [We distributed] y=1/3x+22/3 [We did 6/1=18/3 So 18/3+4/3=22/3] Now we find our equation for this prependicular line and that is y=1/3x+22/3

OpenStudy (anonymous):

thankx now i have anotha one